unseen:

Jimmy has seven times as many dimes as nickels. Their value is $3.75. How many nickels does he have? 35 15 5

unseen:

@dude

unseen:

@tigerlover

tigerlover:

.05n + .10d = 3.75 d = 7n now we sub .05n + .10(7n) = 3.75 .05n + .70n = 3.75 .75n = 3.75 n= 3.75/0.75 n = 5 d = 7n d = 7(5) d = 35 so he has 5 nickels and 35 dimes

unseen:

ok thx